A 12.0 L tank contains 0.305 kg of helium at 16.0 ∘C. The molar mass of helium is 4.00g/mol.

1)How many moles of helium are in the tank?

2)What is the pressure in the tank (in pascals)?

3) What is the pressure in the tank ( in atm)?
Answer #1

1)
Number of moles, n = mass / molecular weight
n = 305/4
n = 76.25 moles

2)
V = 12.0 * 10^-3 m^3
T = 16.0 + 273 = 289 K

Using ideal gas equation
P*V = n*R*T
P*12.0 * 10^-3 = 76.25 * 8.314 * 289
P = 1.527 * 10^7 Pa

(3)
1 atm = 101325 Pa
Pressure in atm, P = 150.67 atm